OpenStudy (anonymous):

Solve: (x+5)/(x^2+x)=1/(x^2+x)-(x-6)/(x+1)

OpenStudy (anonymous):

@jim_thompson5910 @dpaInc @Mertsj

OpenStudy (mertsj):

\[\frac{x+5}{x(x+1)}=\frac{1}{x(x+1)}-\frac{x-6}{x+1}\]

OpenStudy (mertsj):

Common denominator is x(x+1). Clear fractions by multiplying by it:

OpenStudy (mertsj):

\[x+5=1-x(x-6)\]

OpenStudy (mertsj):

\[x+5=1-x^2+6x\] \[x^2-5x+4=0\] \[(x-4)(x-1)=0\] x=4 or x=1

OpenStudy (anonymous):

Thank you so much! =)

OpenStudy (mertsj):

yw